OpenStudy (anonymous):

A cellular Phone company offers a contract for which the cost C, in dollars, of T minutes of telephoning is given by C=0.25(t-600)+75.95, where it is assumed that t > or equal to 600 minutes. What times will keep costs between $120.20 and $147.45?

OpenStudy (anonymous):

C=.25(t-600)+75.95=.25t-150+75.95=.25t-74.05 Because this equation is linear (Read: simple), you can just plug in the two C values (120.20 and 147.45) and get the corresponding t values. 120.20=.25t-75.05, and 147.45=.25t-75.05

OpenStudy (anonymous):

how do i solve 120.20=.25t-75.05, and 147.45=.25t-75.05